# MetaOOD: Automatic Selection of OOD Detection Models

### v1

@@ 3.4 Online Model Selection
Specifically, for a new dataset pair, we acquire the predicted relative performance ranking of different OOD detection methods using the trained $f$, and select the top-1 method [^3] It may choose the top-$k$ candidates as an ensemble, while this work focuses on top-1 selection., as shown in Eq. (2). It is important to note that this procedure is zero-shot, and does not require any network training on the test sample. The (online) model selection steps are given in Fig. 1, bottom and Appx. Algo. 2.
@@ 4 Experiments
Our experiments answer the following research questions (RQ): RQ14.3): How effective is the proposed MetaOOD in unsupervised OOD model selection in comparison to other leading baselines? RQ24.4.1): How do different design choices in MetaOOD impact its effectiveness? RQ34.4.2): How much time overhead/saving MetaOOD brings to OOD detection in general?
@@ 4.1 Experiment Setting
The model Set $\mathcal{M}$. We compose a model set $\mathcal{M}$ to choose from with 11 popular OOD detection models as shown in Table  1, covering different types of detection methods.
@@
Evaluation. To compare the performance of MetaOOD and baselines, we examine the performance rank  [^4] In this work, we choose Area Under the Receiver Operating Characteristic Curve (AUC-ROC), or ROC, as the performance metric; can replace by any other metrics at interest. of the OOD detector chosen by each method among all the candidate detectors through a boxplot and the rank diagram (which is the average across all dataset pairs). Clearly, the best rank is 1, and the worst is 12 (i.e., 11 baselines and MetaOOD) To compare our algorithm with a baseline, we employ the pairwise Wilcoxon rank test on performances across dataset pairs (significance level $p<0.05$). The full selection results are in Appendix Table B.
@@ 4.2 Model Selection Baselines
(a) No model selection or random selection: always employs either the ensemble of all the models or the same single model, or randomly selects a model: (1) Maximum Softmax Probability (MSP) (Hendrycks & Gimpel, 2017), as a popular OOD model, uses the maximum softmax score of a neural network’s output as threshold to identify whether an input belongs to the distribution the network was trained on. (2) ODIN (Liang et al., 2017) applies temperature scaling and small perturbations to the input data, which helps to amplify the difference in softmax scores between ID and OOD samples. (3) Mega Ensemble (ME) averages OOD scores from all the models for a given dataset. As such, ME does not perform model selection but rather uses all the models. (4) Random Selection (Random) randomly selects a model from the pool of candidate models.
@@
(d) Large language models (LLMs) as a model selector: (11) GPT-4o mini (Reimers & Gurevych, 2019) is used as a zero-shot meta-selector. The dataset and method descriptions are directly provided to the LLM, allowing it to select the methods based on these descriptions. Note there is no meta-learning here. The details are presented in Appendix §B.1.
@@ 4.3 Overall Results
First, we present the rank distribution of the actual rank of the top-1 OOD detector chosen by each model selection method across the 24 test data pairs in Fig. 3. To compare two model selection algorithms (e.g., ours with a baseline), we perform Wilcoxon rank test on the AUROC value of the top-1 models selected by our method and the baseline method, as shown in Table 2. Also, we show the aggregated average rank plot in Fig. 2, where the average performance rank of the OOD detection model selected by each algorithm is shown. Here are the key findings of the results:
1. MetaOOD outperforms all baselines, as shown in Fig. 3, Table 2, and Fig. 2. As shown in Fig. 3, MetaOOD consistently performs well with small variance. On the aggregated level, the average rank of MetaOOD surpasses all of the 11 baselines in Fig. 2. Moreover, MetaOOD has demonstrated statistically superior performance compared to all these baselines. Such enhancement in performance is indicative of MetaOOD’s robust approach to tackling complex datasets while generating stable performance. We credit its superiority to the combination of a meta-learning framework and the use of language models in embedding datasets and models.
2. Optimization-based meta-learning methods (i.e., MetaOOD, MetaOOD_0 and ALORS) perform better in unsupervised model selection over other baselines. One reason is that meta-learning leverages prior experiences to adapt to new tasks. It optimizes the learning process by extracting common patterns and representations across different tasks, enhancing the model’s generalization ability. Meanwhile, the optimization process helps models converge to an optimal solution efficiently. Compared to simple meta-learners, such optimization-based methods take advantage of the meta-features for a better mapping of the model performance. For instance, AS and ISAC do not involve optimization and resemble finding the closest sample or closest sample cluster for choosing the top-1 detector. They are inferior to MetaOOD as there is no optimization on the OOD model performance, but fully depends on the data embeddings for performance knowledge transfer.
@@
  • ME: Averaging the OOD scores of all models does not yield strong results, as demonstrated in Fig. 3 and 2. This may stem from certain models consistently underperform across various datasets, and combining all models indiscriminately reduces overall effectiveness. While using selective ensembles might offer improvements (Zhao & Hryniewicki, 2019), constructing ensembles from numerous models can be impractical due to high costs. In contrast, MetaOOD learns to make optimal selections without constructing any models, allowing it to operate efficiently during testing.
  • Random: According to Fig. 3 and Table 2, random selection performs worse than all the meta-learners. This indicates that all the meta-learner baselines we chose do have some improvements compared to random choice, and it is not advised to select an OOD detection model randomly.
@@
4. LLM acts as a reasonable zero-shot model selector in OOD model selection. This highlights the potential of LLM in model selection, while it can be improved by meta-learning. While the performance of LLM as a selector may be slightly worse (ranked as top 3) than our MetaOOD, LLM consistently selects the globally best model (Openmax (Bendale & Boult, 2016)) across diverse tasks, with no prior knowledge about the performances of different OOD detection methods on different models. This is particularly valuable when prior knowledge is limited or when facing novel tasks, highlighting the potential of LLMs in unsupervised model selection. It can be further improved via meta-learning by receiving more performance-related information. Arguably, MetaOOD leverages language embeddings along with the meta-learning, thus achieving better performance.
@@ 4.4.1 Ablation Study
  • LLama2-7b (Llama) (Touvron et al., 2023): LLama2 is an advanced language model that provides enhanced capabilities in a variety of applications, including text summarization, translation, and conversational AI. Since it is a transformer style decoder-based model, the sequence level embeddings are produced by pooling token level embeddings together [^5] Usually this is done by averaging the token level embeddings or using the first token. In this study, we opt for the latter approach..
@@ 5 Conclusion, Limitations, and Future Directions
Ethics Statement: Our research adheres to the general Code of Ethics, ensuring that MetaOOD is developed and applied with ethical considerations at the forefront, particularly in privacy, bias, and fairness across diverse applications. By facilitating more accurate and unbiased model selection, MetaOOD helps mitigate potential ethical risks in its deployments, such as in surveillance and healthcare, promoting fairness and protecting privacy. Continuous ethical evaluations accompany MetaOOD’s development to ensure it meets societal and legal standards.
Reproducibility Statement: We advocate reproducibility in MetaOOD. Comprehensive documentation of our methodologies and experimental designs is detailed in the main text and appendices. We have made our code, testbed, and meta-learner fully accessible at https://anonymous.4open.science/r/metaood-DA67, providing the resources for replication and exploration.
@@ A.1 Pseudo-code for Meta-train and Online Model Selection
A.2 Details on Meta-features
The classical way of data embedding via meta-features is discussed in §3.3.1. Table A presents the complete list of meta-features constructed in this study.
@@ A.2 Details on Meta-features
Table A: Selected MetaOOD features. Part of the statistical features are based on Vanschoren (2018); Zhao et al. (2021), and specialized landmarker features are newly designed for OOD detection. See §3.3 for more details.
@@ B.2 Full Performance Results
Below is the full performance matrix $P$, which shows the performance of eleven OOD detection models (refer to Table 1) on our constructed ID-OOD dataset pairs (see dataset pair construction in §4.1). We report AUROC as the metric for OOD detection.
@@
Table B: Various OOD detection models’ performance on ID-OOD dataset pairs. See experiment setting in §4.1.

### v2

@@ 3.4 Online Model Selection
Specifically, for a new dataset pair, we acquire the predicted relative performance ranking of different OOD detection methods using the trained $f$, and select the top-1 method [^3] It may choose the top-$k$ candidates as an ensemble, while this work focuses on top-1 selection., as shown in Eq. (2). It is important to note that this procedure is zero-shot, and does not require any training on the test sample. The (online) model selection steps are given in Fig. 1, bottom and Appx. Algo. 2.
@@ 4 Experiments
Our experiments answer the following research questions (RQ): RQ14.3): How effective is the proposed MetaOOD in unsupervised OOD detection model selection in comparison to other leading baselines? RQ24.4.1): How do different design choices in MetaOOD impact its effectiveness? RQ34.4.2): How much time overhead/saving MetaOOD brings to OOD detection in general?
@@ 4.1 Experiment Setting
The model Set $\mathcal{M}$. We compose a model set $\mathcal{M}$ to choose from with 11 popular OOD detection models as shown in Tab. 1, covering different types of detection methods.
@@
Evaluation. To compare the performance of MetaOOD and baselines, we examine the performance rank [^4] In this work, we choose Area Under the Receiver Operating Characteristic Curve (AUC-ROC), or ROC, as the performance metric; can replace by any other metrics at interest. of the OOD detector chosen by each method among all the candidate detectors through a boxplot and the rank diagram (which is the average across all dataset pairs). Clearly, the best rank is 1, and the worst is 12 (i.e., 11 baselines and MetaOOD). To compare our algorithm with a baseline, we employ the pairwise Wilcoxon rank test on performances across dataset pairs (significance level $p<0.05$). The full selection results are in Appx. Tab. C.
@@ 4.2 Model Selection Baselines
(a) No model selection or random selection: always employs either the ensemble of all the models or the same single model, or randomly selects a model: (1) Maximum Softmax Probability (MSP) (Hendrycks & Gimpel, 2017), as a popular OOD detection model, uses the maximum softmax score of a neural network’s output as threshold to identify whether an input belongs to the distribution the network was trained on. (2) ODIN (Liang et al., 2017) applies temperature scaling and small perturbations to the input data, which helps to amplify the difference in softmax scores between ID and OOD samples. (3) Mega Ensemble (ME) averages OOD scores from all the models for a given dataset. As such, ME does not perform model selection but rather uses all the models. (4) Random Selection (Random) randomly selects a model from the pool of candidate models.
@@
(d) Large language models (LLMs) as a model selector: (11) GPT-4o mini (OpenAI et al., 2024) is used as a zero-shot meta-selector. The dataset and method descriptions are directly provided to the LLM, allowing it to select the methods based on these descriptions. Note there is no meta-learning here. The details are presented in Appx. §B.1.
@@ 4.3 Overall Results
First, we present the rank distribution of the actual rank of the top-1 OOD detector chosen by each model selection method across the 24 test data pairs in Fig. 3. To compare two model selection algorithms (e.g., ours with a baseline), we perform Wilcoxon rank test on the AUROC value of the top-1 models selected by our method and the baseline method, as shown in Tab. 2. Also, we show the aggregated average rank plot in Fig. 2, where the average performance rank of the OOD detection model selected by each algorithm is shown. Here are the key findings of the results:
1. MetaOOD outperforms all baselines. As shown in Fig. 3, MetaOOD consistently performs well with small variance. On the aggregated level, the average rank of MetaOOD surpasses all of the 11 baselines in Fig. 2. Moreover, MetaOOD has demonstrated statistically superior performance compared to all these baselines, as shown in Tab. 2. Such enhancement in performance is indicative of MetaOOD’s robust approach to tackling complex datasets while generating stable performance. We credit its superiority to the combination of a meta-learning framework and the use of language models in embedding datasets and models.
2. Optimization-based meta-learning methods (i.e., MetaOOD, MetaOOD_0 and ALORS) perform better in unsupervised model selection over other baselines. One reason is that meta-learning leverages prior experiences to adapt to new tasks. It optimizes the learning process by extracting common patterns and representations across different tasks, enhancing the model’s generalization ability. Meanwhile, the optimization process helps models converge to an optimal solution efficiently. Compared to simple meta-learners, such optimization-based methods take advantage of the meta-features for a better mapping of the model performance. For instance, AS and ISAC do not involve optimization and resemble finding the closest sample or closest sample cluster for choosing the top-1 detector. They are inferior to MetaOOD as there is no optimization on the OOD detector selection, but fully depends on the data embeddings for performance knowledge transfer.
@@
  • ME: Averaging the OOD detection scores of all models does not yield strong results, as demonstrated in Fig. 2 and 3. This may stem from certain models consistently underperform across various datasets, and combining all models indiscriminately reduces overall effectiveness. While using selective ensembles might offer improvements (Zhao & Hryniewicki, 2019), constructing ensembles from numerous models can be impractical due to high costs. In contrast, MetaOOD learns to make optimal selections without constructing any models, allowing it to operate efficiently during testing.
  • Random: According to Fig. 3 and Tab. 2, random selection performs worse than all the meta-learners. This indicates that all the meta-learner baselines we chose do have some improvements compared to random choice, and it is not advised to select an OOD detection model randomly.
@@
4. LLM acts as a reasonable zero-shot model selector in OOD detection model selection. This highlights the potential of LLM in model selection, while it can be improved by meta-learning. While the performance of LLM as a selector may be slightly worse (ranked as top 3) than our MetaOOD, LLM consistently selects the globally best model (Openmax (Bendale & Boult, 2016)) across diverse tasks, with no prior knowledge about the performances of different OOD detection methods on different models. This is valuable when prior knowledge is limited or when facing novel tasks, highlighting the potential of LLMs in unsupervised model selection. It can be improved via meta-learning by receiving more performance-related information. Arguably, MetaOOD leverages language embeddings along with the meta-learning, thus achieving better performance. Future work may consider LLMs’ robustness and trustworthiness in model selection Huang et al. (2025).
@@ 4.4.1 Ablation Study
  • LLama2-7b (Llama) (Touvron et al., 2023): LLama2 is an advanced language model that provides enhanced capabilities in a variety of applications, including text summarization, translation, and conversational AI. Since it is a transformer style decoder-based model, the sequence level embeddings are produced by pooling token level embeddings together [^5] Usually this is done by averaging the token level embeddings or using the last token. In this study, we opt for the latter approach..
@@ 5 Conclusion, Limitations, and Future Directions
Ethics Statement: Our research adheres to the ICLR Code of Ethics, ensuring that MetaOOD is developed and applied with ethical considerations at the forefront, particularly in privacy, bias, and fairness across diverse applications. By facilitating more accurate and unbiased model selection, MetaOOD helps mitigate potential ethical risks in its deployments, such as in surveillance and healthcare, promoting fairness and protecting privacy. Continuous ethical evaluations accompany MetaOOD’s development to ensure it meets societal and legal standards.
Reproducibility Statement: We advocate reproducibility in MetaOOD. Comprehensive documentation of our methodologies and experimental designs is detailed in the main text and appendices. We have made our code, testbed, and meta-learner fully accessible at https://github.com/yqin43/metaood, providing the resources for replication and exploration.
Acknowledgments
This work was partially supported by the National Science Foundation under Award No. 2346158. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.
@@ A.1 Pseudo-code for Meta-train and Online Model Selection
A.2 Details on Notations
The following notations are used in Fig. 1, which provides a comprehensive MetaOOD overview (§3.2).
Notations Description
$\mathcal{L}$ Training Loss
$\mathcal{M}$ OOD Detection Methods
$\mathcal{N}$ Dataset Pairs
$\phi$ Embedding Notation for OOD Detection Methods
$\psi$ Embedding Notation for Dataset Pairs
$P_{i,j}$ Performance of OOD Detection Method $j$ on Dataset Pair $i$
$\hat{P}_{i,j}$ Predicted Performance of OOD Detection Method $j$ on Dataset Pair $i$
$f$ Meta-predictor
$X$ Input Sample
Table A: Notations with details used in Fig. 1
A.3 Details on Meta-features
The classical way of data embedding via meta-features is discussed in §3.3.1. Tab. B presents the complete list of meta-features constructed in this study.
@@ A.3 Details on Meta-features
Table B: Selected MetaOOD features. Part of the statistical features are based on Vanschoren (2018); Zhao et al. (2021), and specialized landmarker features are newly designed for OOD detection. See §3.3 for more details.
@@ B.2 Full Performance Results
Below is the full performance matrix $P$, which shows the performance of eleven OOD detection models (refer to Tab. 1) on our constructed ID-OOD dataset pairs (see dataset pair construction in §4.1). We report AUROC as the metric for OOD detection.
@@
Table C: Various OOD detection models’ performance on ID-OOD dataset pairs. See experiment setting in §4.1. We highlight the selected OOD method for each dataset on the test set in bold.
B.3 Dataset Descriptions
Below is the full list of the dataset descriptions we use in the study, which includes basic dataset metadata such as the types of objects in each dataset, the image type, and the size of the dataset. This information can be compiled swiftly and easily.
Dataset Description
CIFAR-10 Contains images of 10 types of objects, including airplanes, cars, birds, cats, deer, dogs, frogs, horses, ships, and trucks. Each image is a small 32x32 RGB image.
CIFAR-100 Similar to CIFAR-10 but includes 100 classes grouped into 20 superclasses. Each class contains images of specific objects, animals, or people.
MNIST Comprises black-and-white images of handwritten digits from 0 to 9, with each image normalized to fit in a 28x28 pixel bounding box.
Places365 Features images of various places like homes, parks, offices, and cities, spread across 365 different scene categories.
SVHN Includes images of digit sequences found in natural scene images, primarily focusing on house numbers from street view images.
Texture Generally consists of images depicting various surface textures such as bricks, tiles, fabrics, and other patterned surfaces.
TIN Contains tiny 32x32 images sourced from the internet, depicting a wide variety of everyday objects, animals, and scenes.
SSB-Hard Contains images of specific breeds of birds, types of aircraft, and models of cars.
NINCO Contains natural scenes, textures, and objects, providing a diverse set of images.
iNaturalist Consists of images submitted by users of various species of plants, animals, and fungi, used for species identification and classification.
Textures Focuses on images that capture the surface quality of various materials.
OpenImages Features a wide range of images with various objects, scenes, and activities, annotated with labels and bounding boxes.
ImageNet Contains diverse images across a wide range of categories like different types of animals, plants, vehicles, and everyday objects.
Fashion-MNIST Includes images of fashion products from 10 categories, such as shirts, dresses, shoes, and bags. Each image is a 28x28 grayscale image.
Table D: Dataset descriptions used in the study, which contain basic information of the dataset such as dataset content (e.g., what kind of objects are in the dataset), image type, and dataset size.
B.4 Feature Importance of Language Embeddings
Fig. A shows the feature importance of the language embeddings, where the F score of a feature is the total number of times the feature is used to split the data in all trees in the model. We show the top-10 most important features. According to the analysis, the initial dimensions of these embeddings carry greater significance.
Tab. E presents the average feature importance based on three metrics: 1) Weight: the number of times a feature is used to split the data across all trees, 2) Cover: the average coverage across all splits the feature is used in, and 3) Gain: the average gain across all splits the feature is used in the ID dataset embeddings. These metrics are computed for the ID dataset embeddings, OOD dataset embeddings, and method embeddings. The results indicate that dataset embeddings have a greater impact on the selection process compared to method embeddings.
Figure A: Top-10 feature importance of the language embeddings used in the study; dimensions from the ID dataset are labeled as ID_dim, those from the OOD dataset as OOD_dim, and from the Method as Method_dim. The initial dimensions of language embeddings are the most critical.
Metric Embeddings ID dataset OOD dataset Method
Weight 0.53 0.52 0.37
Cover 7.27e-07 2.88e-07 7.08e-07
Gain 0.026 0.025 0.018
Table E: Average feature importance for different language embeddings on the Weight, Cover, and Gain metrics. Dataset embeddings play a more significant role in the selection process compared to method embeddings.
Appendix C Further study on embeddings
C.1 Combined meta-feature and language embedding
Some might argue that relying solely on language embeddings could be overly basic. To address this, we also experimented with combining statistical and landmarker meta-features alongside language embeddings for dataset and method embeddings (denoted as Combined). As shown in Fig. B, this approach achieves performance comparable to MetaOOD. However, the computational and time costs of extracting statistical features can be substantial, particularly for large-scale datasets like ImageNet and LSUN. Language embeddings were chosen primarily for their reproducibility, efficiency, and significantly lower computational overhead.
C.2 Variation of Dataset Descriptions
To test the robustness of language embeddings against different input descriptions of datasets, we manually adjust dataset descriptions by rephrasing them to alter the wording while preserving the original meaning, as detailed in Tab. F. This variation is referred to as MetaOOD’. According to Fig. C, changes in dataset descriptions containing basic information do not lead to significant differences, and performance remains stable.
Figure B: Boxplot of the rank distribution of combined meta-feature and language embedding, and baselines. The performance is comparable to that of MetaOOD.
Figure C: Boxplot of the rank distribution of variation of dataset descriptions and baselines. Variations in dataset descriptions, which include basic information, do not lead to significant differences, and performance remains stable.
Dataset Description
CIFAR-10 A standard dataset in computer vision, containing 60,000 color images (32x32 pixels) of 10 distinct object categories: airplanes, cars, birds, cats, deer, dogs, frogs, horses, ships, and trucks.
CIFAR-100 An extension of CIFAR-10, featuring 100 classes grouped into 20 broader superclasses. It includes 60,000 color images (32x32 pixels), each class representing a finer granularity of objects, animals, or people.
MNIST A seminal dataset comprising 70,000 grayscale images (28x28 pixels) of handwritten digits (0-9), each centered in a bounding box.
Places365 Contains 1.8 million images spanning 365 distinct scene categories, including various indoor and outdoor environments like homes, parks, offices, and cities.
SVHN A real-world image dataset featuring over 600,000 images of digit sequences (32x32 pixels) extracted from house numbers in street view imagery.
Texture Composed of diverse images highlighting surface textures such as bricks, tiles, fabrics, and other patterned materials, often used to study material properties.
TiN A massive dataset of 80 million tiny images (32x32 pixels) collected from the web, representing a broad spectrum of everyday objects, animals, and scenes.
SSB Hard Features a diverse collection of images, including specific bird species, distinct types of aircraft, and detailed models of cars.
NINCO Comprises a wide range of images, including natural scenes, intricate textures, and various objects.
iNaturalist Comprises over 859,000 images of various species of plants, animals, and fungi, contributed by users for species identification and classification.
Textures Similar to the Texture dataset, but may have a different focus or a wider range of surface textures, including natural and artificial materials.
OpenImages A vast dataset containing 9 million images annotated with labels, bounding boxes, and object segmentation across diverse objects, scenes, and activities.
ImageNet One of the most comprehensive datasets, with over 14 million images categorized into 1,000 classes, including animals, plants, vehicles, and various everyday objects.
Fashion-MNIST A modern alternative to MNIST, containing 70,000 grayscale images (28x28 pixels) of fashion items categorized into 10 classes, such as shirts, dresses, shoes, and bags.
Table F: Manually modified dataset descriptions used in additional study C.2, which contain basic information of the dataset such as dataset content (e.g., what kind of objects are in the dataset), image type, and dataset size.